OpenStudy (anonymous):

If customer A has an appointment 2 2/3 hours before customer B who has an appointment 5/6 of an hour before customer C. How much later is customer C‘s appointment than customer A’s? An answer is requiredis it 3.5 hrd

OpenStudy (anonymous):

A + 2 + 2/3 = B B + 5/6 = C so A + 2 + 2/3 + 5/6 = C C - A = ?